1.IF THE SUM OF TEN CONSECUTIVE POSITIVE INTEGERS IS 115, FIND THE LARGEST OF THESE INTEGERS.
2.THREE CONSECUTIVE INTEGERS HAVE A SUM OF 12 AND A PRODUCT OF 60. IF THE SMALLEST INTEGER IS 3 FIND THE LARGEST NUMBER.

Answer by chana(69)

1.IF THE SUM OF TEN CONSECUTIVE POSITIVE INTEGERS IS 115, FIND THE LARGEST OF THESE INTEGERS.
sol :
let the nos be x , x+1, x+2, x+3, x+4, x+5, x+6, x+7, x+8, x+9
so 10x + 45 = 115
solve tis you will get x = 7
substitute in the 10th number = x + 9
largest no = 16
2.THREE CONSECUTIVE INTEGERS HAVE A SUM OF 12 AND A PRODUCT OF 60. IF THE SMALLEST INTEGER IS 3 FIND THE LARGEST NUMBER.
sol :
if first no is 3 so ,the third consecutive number is 3 + 2 = 5 , which is the largest.